# Integrated Positioning and Communication via LEO Satellites: Opportunities and Challenges

### v1

@@ I Introduction
With the rise of applications such as the internet of everything, self-driving vehicles, and spatial intelligence, the demand for accurate positioning and reliable communication continues to grow. As a pivotal technology, low Earth orbit (LEO) satellites nowadays are playing an important role in the fields of mobile radio positioning and communication in fifth generation (5G) and beyond. In this section, we begin by introducing the unique features of LEO satellite systems, followed by a brief review of their positioning and communication capabilities.
@@ I-A LEO Features
leo satellites are positioned at orbits altitudes ranging from 160 km to 2,000 km. Compared to medium Earth orbit (MEO) and geostationary orbit (GEO) satellites, their lower orbital altitude significantly reduces the signal propagation distance, which results in lower path loss and thus stronger received signals, as well as reduced delay. Furthermore, the wide distribution and large number of LEO satellites enable global coverage and consistent visibility to ground users. Notable constellations such as Starlink, OneWeb, and Telesat, which comprise hundreds to thousands of satellites, form effective positioning and communication networks. Notably, recent technological advancements have also significantly reduced the costs associated with the satellites’ fabrication and launching. The advancement of smaller, lighter microsatellites, requiring fewer materials, has facilitated the use of more cost-effective launch vehicles. Additionally, the utilization of Reusable Launch Vehicles (RLVs) has further reduced costs and improved the efficiency of space launches. The quick turnaround ability of RLVs also allows for a substantial increase in launch frequency [1].
@@ I-A2 Space-Based Network Architecture
Unlike the ground-based network where satellites simply relay signals, the space-based network architecture enables more complex communication tasks. The space segment often utilizes regenerative payloads, which not only perform the functions of transparent payloads but also understand and process signals, such as coding/decoding and modulation/demodulation [3]. Additionally, there exist inter-satellite links connecting these regenerative payloads, thereby enabling collaborative designs among the satellites with regard to both positioning and communication.
@@ I-B LEO Satellite-Based Positioning
In recent years, positioning based on LEO satellites, which forms a part of the broader framework of LEO satellite positioning, navigation, and timing (PNT), has attracted significant attention for its potential to complement global navigation satellite system (GNSS) or serve as a reliable backup in case GNSS systems fail. LEO PNT can be implemented through several strategies [4]. One effective strategy employs dedicated LEO satellites, which are specifically engineered to deliver optimal PNT services. Alternatively, LEO satellites not initially designed for PNT can still be utilized opportunistically to gather PNT data through signals of opportunity (SoOP). In such systems, LEO satellite-based communication systems can be enhanced by PNT services. This enhancement can be achieved by either integrating a PNT-specific payload into existing communication satellites or by incorporating PNT functionalities directly into the physical layer of satellite communication systems.
Positioning with LEO satellites provides global coverage and offers substantial advantages over systems that utilize satellites in higher orbits, such as MEO and GEO. The proximity of LEO satellites to Earth results in higher received signal strength, which not only enhances positioning accuracy but also facilitates effective indoor signal penetration where GNSS fails to reach. This robust signal strength also bolsters security by offering strong resistance to interference, including jamming and spoofing attacks. Additionally, the fast-moving nature of LEO satellites allows the utilization of Doppler shift for precise positioning [4]. Moreover, the significantly larger constellations of LEO satellites, outnumbering GNSS constellations by several orders of magnitude, have the potential to offer higher positioning accuracy through increased observable redundancy and expanded coverage, particularly in urban environments with open-sky conditions.
@@ I-C LEO Satellite-Based Communication
The main differences between LEO communication and terrestrial communication lie in the severer path loss, the greater propagation delay, and the stronger Doppler effect.
@@
Attenuation due to either ionospheric or tropospheric scintillation Corresponds to rapid fluctuations of the received signal amplitude and phase. Scintillations depend on location, time-of-day, season, solar and geomagnetic activity. Ionosphere scintillation shall only be considered for frequencies below 6 GHz, while tropospheric propagation shall only be considered for frequencies above 6 GHz.
@@
The path loss model of LEO satellite channel typically consists of two components: the large-scale path loss and the small-scale fading. According to 3GPP TR 38.811 [3], the large-scale path loss of satellite-terrestrial channels comprises several components: free space path loss, shadow fading loss, clutter loss, atmospheric absorption, ionospheric/tropospheric scintillation, and potential building penetration loss. A summary of these attenuation effects and their key characteristics is listed in Table I. In addition to large-scale path loss, small-scale fading, or fast fading, must be considered. Most satellite channel models rely on flat fading. The statistical model outlined in [6] for mixed propagation conditions in the land mobile-satellite service is one of the most up-to-date references. Detailed modeling methodology of frequency-selective fading models is described in Section 7.5 of 3GPP TR38.901[7] and Section 6.7.2 of 3GPP TR38.811[3]. These models indicate that the path loss in satellite-terrestrial links is significantly higher than in terrestrial networks due to not only longer propagation distances but also a complex atmospheric environment.
The long propagation distance of satellite channels also results in substantial propagation delays, which increases communication latency. Additionally, compared to GNSS, the lower orbital altitude of LEO satellites causes more significant variations in propagation distance to different ground users, leading to highly diverse propagation delays from different users. Furthermore, considering the fast movement of LEO satellites, these propagation delays are also time-varying. This effect can potentially hinder the usability of delay-sensitive services. Consequently, some delay-related metrics like delay satisfaction ratio and delay variance are adopted as key performance indicators of the overall LEO satellite network.
@@ II-A1 Position Information-Enabled Fast Beamforming
With the ability to concentrate transmit power in targeted directions, multi-antenna beamforming is crucial in combating severe path loss and enhancing LEO satellite communication capacity, just as it does in terrestrial networks. However, beamforming techniques typically require channel state information (CSI). As analyzed in the previous subsection, LEO satellite channels are facing the unique challenge of short coherence time and fast variation. This imposes significant challenges for real-time channel estimation and beamforming. Considering the LoS-dominant propagation and the known satellite orbits, CSI-free beamforming that leverages positioning information represents a potential solution. Given the locations of the users and the motion of the satellites, directional beamformers can be designed predictively and dynamically towards the LoS direction. This approach holds great promise for avoiding the high time cost associated with channel estimation. However, beamforming based on the LoS channel only may result in a lack of diversity, which can undermine the reliability of LEO satellite communication systems. With the increasingly dense deployment of LEO satellites, multi-satellite-enabled cooperative beamforming, which potentially enriches diversity and boosts received power, can be a promising solution.
@@ II-B1 Infrastructure Perspective
Owing to using higher frequency bands, LEO satellite communication systems can employ advanced infrastructures such as phased array antennas and hybrid beamforming structures to enhance communication performance, which also benefit positioning applications significantly. For instance, these infrastructures enable the acquisition of the angle-of-departure (AoD) and angle-of-arrival (AoA) observations, which are unavailable in GNSS that uses single-antenna transmitters and receivers. These angle observations are closely related to the UE position and can thus be explored to boost positioning performance. Moreover, when massive multiple-input multiple-output (MIMO) technology is deployed on LEO satellites, the accuracy of these angle observations can be further improved because of the higher spatial resolution. Besides, in 5G/ sixth generation (6G) systems, the extensive terrestrial communication infrastructure can be leveraged and integrated with satellite-based positioning systems to enhance localization capabilities. Traditional GNSS applications already demonstrate the use of ground stations to assist high-precision positioning, as seen in real-time kinematic (RTK) positioning systems [11]. In emerging LEO satellite systems, theoretical studies have also shown that integrating ground-based technologies, such as reconfigurable intelligent surface (RIS), have the potential to significantly enhance positioning and tracking accuracy, particularly in complex suburban and urban environments [12].
@@ II-B2 Network Coordination Perspective
In addition to the advantages provided by new communication infrastructures, efficient inter-satellite and satellite-to-ground communication can also significantly enhance positioning performance by enabling cooperation across the network. For example, multiple LEO satellites equipped with MIMO antenna arrays can achieve superior performance through coordinated beamforming, as opposed to independent beamforming schemes. However, this coordination is only feasible with efficient inter-satellite communication links. Furthermore, it is preferable for the signals from each satellite to be orthogonal (in time or frequency), to mitigate mutual interference and ensure accurate positioning, which requires communication and coordination between satellites. Therefore, information sharing between different positioning anchors (satellites or BSs), enabled by efficient communication links, is the key to enhancing the positioning performance of the entire system.
@@ II-B3 Other Benefits
LEO communication technologies enable satellites to synchronize their clocks with ground-based atomic clocks via signals, ensuring highly accurate time data for position calculations. Additionally, LEO communication technologies facilitate the implementation of fault-tolerant systems capable of automatically switching control and rerouting data in the event of satellite or ground station failures. This redundancy is essential for maintaining uninterrupted positioning services. Furthermore, the two-way communication capability of LEO systems enhances dynamic and responsive positioning services, providing an improvement over systems lacking a return channel.
@@ III-A Simulation Setup
Considering a downlink scenario, we simulate LEO satellites at an altitude of 400 km, each with a transmit power of 60 dBm and a 240 MHz bandwidth, operating at 28 GHz. The ground UE is modeled as a single-antenna receiver with a noise power spectral density (PSD) of -174 dBm/Hz. The wireless channels are generated using the Rician fading model, with a Rician factor of 3 by assuming the user is in a rural environment [12, Table II]. By default, the LEO satellites are equipped with a $20\times 20$ antenna array with half-wavelength spacing, oriented towards the Earth. For simplicity, we employ an analog beamforming schemes.
@@ IV-A Doppler Shift
The strong Doppler effect on LEO satellite systems is a double-edged sword. On the one hand, the Doppler shift provides valuable information about the relative velocity between users and satellites, which is beneficial to positioning. On the other hand, this effect can be harmful to communication. As mentioned in Section I-C, an approximate 700 kHz Doppler frequency is comparable to the communication symbol rate. This will lead to severe inter-subcarrier interference and further degrade the demodulation performance. Additionally, the high Doppler shift variation in LEO systems may exceed the tolerances that a communication system can support [14]. Thus, advanced algorithms are required to balance the utilization of the Doppler effect for accurate positioning and the mitigation of its negative effects in communication.
@@ IV-C Satellite Orbit Errors
The orbit information of LEO satellites is typically not accurately known due to several factors. Perturbing forces, such as atmospheric drag and solar radiation pressure, are primary contributors to these inaccuracies. Moreover, the parametrization of the orbit stored in two-line element (TLE) files further complicates precise orbit prediction [4]. These factors introduce orbital errors, causing mismatches in the satellites’ position and orientation information, which ultimately degrade positioning and communication performance. Fig. 5 presents an evaluation of the positioning RMSE versus channel observation accuracy (using channel delay as a representative) under different levels of satellite position mismatch. The results indicate that such mismatches can lead to performance saturation, limiting the achievable accuracy of the system. A similar negative impact can be observed in communication performance, especially in an IPAC system.
@@ VI Acknowledgments
Fig. 1 was created by Ana Runte, Scientific Illustrator at KAUST.

### v2

@@ I Introduction
With the rise of applications such as the internet of everything, self-driving vehicles, and spatial intelligence, the demand for accurate positioning and reliable communication continues to grow. As a pivotal technology, low Earth orbit (LEO) satellites nowadays are playing an important role in the fields of mobile radio positioning and communication in fifth generation (5G) and beyond [1]. In this section, we begin by introducing the unique features of LEO satellite systems, followed by a brief review of their positioning and communication capabilities.
@@ I-A LEO Features
leo satellites are positioned at orbit altitudes ranging from 160 km to 2,000 km. Compared to medium Earth orbit (MEO) and geostationary orbit (GEO) satellites, their lower orbital altitude significantly reduces the signal propagation distance, which results in lower path loss and thus stronger received signals, as well as reduced delay. Furthermore, the wide distribution and large number of LEO satellites enable global coverage and consistent visibility to ground users. Notable constellations such as Starlink, OneWeb, and Telesat, which comprise hundreds to thousands of satellites, form effective positioning and communication networks. Notably, recent technological advancements have also significantly reduced the costs associated with the satellites’ fabrication and launching. The advancement of smaller, lighter microsatellites, requiring fewer materials, has facilitated the use of more cost-effective launch vehicles. Additionally, the utilization of Reusable Launch Vehicles (RLVs) has further reduced costs and improved the efficiency of space launches. The quick turnaround ability of RLVs also allows for a substantial increase in launch frequency [2].
@@ I-A2 Space-Based Network Architecture
Unlike the ground-based network where satellites simply relay signals, the space-based network architecture enables more complex communication tasks. The space segment often utilizes regenerative payloads, which not only perform the functions of transparent payloads but also understand and process signals, such as coding/decoding and modulation/demodulation [4]. Additionally, these regenerative payloads can be interconnected through inter-satellite links, thereby enabling more flexible and advanced system architectures.
@@ I-B LEO Satellite-Based Positioning
In recent years, positioning based on LEO satellites, which forms a part of the broader framework of LEO satellite positioning, navigation, and timing (PNT), has attracted significant attention for its potential to complement global navigation satellite system (GNSS) or serve as a reliable backup in case GNSS systems fail. LEO PNT can be implemented through several strategies [5]. One effective strategy employs dedicated LEO satellites, which are specifically engineered to deliver optimal PNT services. Alternatively, LEO satellites not initially designed for PNT can still be utilized opportunistically to gather PNT data through signals of opportunity (SoOP). In such systems, LEO satellite-based communication systems can be enhanced by PNT services. This enhancement can be achieved by either integrating an additional PNT-specific payload into existing communication satellites or by incorporating PNT functionalities directly into the communication payload [5].
Positioning with LEO satellites provides global coverage and offers substantial advantages over systems that utilize satellites in higher orbits, such as MEO and GEO. The proximity of LEO satellites to Earth results in higher received signal strength, which not only enhances positioning accuracy but also facilitates effective indoor signal penetration where GNSS fails to reach. This robust signal strength also bolsters security by offering strong resistance to interference, including jamming and spoofing attacks. Additionally, the fast-moving nature of LEO satellites allows the utilization of Doppler shift for precise positioning [5]. Moreover, the significantly larger constellations of LEO satellites, outnumbering GNSS constellations by several orders of magnitude, have the potential to offer higher positioning accuracy through increased observable redundancy and expanded coverage, particularly in urban environments with unobstructed sky views.
@@ I-C LEO Satellite-Based Communication
The main differences between LEO communication and terrestrial communication lie in the more severe path loss, the greater propagation delay, and the stronger Doppler effect.
@@
Attenuation due to either ionospheric or tropospheric scintillation Corresponds to rapid fluctuations of the received signal amplitude and phase. Scintillations depend on location, time-of-day, season, solar and geomagnetic activity. Ionosphere scintillation shall only be considered for frequencies below 6 GHz, while tropospheric scintillation shall only be considered for frequencies above 6 GHz.
@@
The path loss model of LEO satellite channel typically consists of two components: the large-scale path loss and the small-scale fading. According to 3GPP TR 38.811 [4], the large-scale path loss of satellite-terrestrial channels comprises several components: free space path loss, shadow fading loss, clutter loss, atmospheric absorption, ionospheric/tropospheric scintillation, and potential building penetration loss. A summary of these attenuation effects and their key characteristics is listed in Table I. In addition to large-scale path loss, small-scale fading, or fast fading, must be considered. Most satellite channel models are based on flat fading. The statistical model outlined in [7] for mixed propagation conditions in the land mobile-satellite service is one of the most up-to-date references. Detailed modeling methodology of frequency-selective fading models is described in Section 7.5 of 3GPP TR38.901[8] and Section 6.7.2 of 3GPP TR38.811[4]. These models indicate that the path loss in satellite-terrestrial links is significantly higher than in terrestrial networks due to not only longer propagation distances but also a complex atmospheric environment.
The long propagation distance of satellite channels also results in substantial propagation delays, which increases communication latency. Additionally, compared to GNSS, the lower orbital altitude of LEO satellites causes more significant variations in propagation distance to different ground users, leading to highly diverse propagation delays to different users. Furthermore, considering the fast movement of LEO satellites, these propagation delays are also time-varying. This effect can potentially hinder the usability of delay-sensitive services. Consequently, some delay-related metrics like delay satisfaction ratio and delay variance are adopted as key performance indicators of the overall LEO satellite network.
@@ II-A1 Position Information-Enabled Fast Beamforming
With the ability to concentrate transmit power in targeted directions, multi-antenna beamforming is crucial in combating severe path loss and enhancing LEO satellite communication capacity, just as it does in terrestrial networks. However, beamforming techniques typically require channel state information (CSI). As analyzed in the previous subsection, LEO satellite channels are facing the unique challenge of short coherence time and fast variation. This imposes significant challenges for real-time channel estimation and beamforming. Considering the LoS-dominant propagation and the known satellite orbits, CSI-free beamforming that leverages positioning information represents a potential solution. Given the locations of the users and the motion of the satellites, directional beamformers can be designed predictively and dynamically towards the LoS direction. This approach holds great promise for avoiding the high time cost associated with channel estimation. However, beamforming for LoS-oriented illumination may inadvertently constrain signal diversity, which is conventionally enriched by multipath propagation. This limitation can potentially undermine the system’s communication reliability, particularly in environments characterized by signal obstructions or complex propagation conditions. With the increasingly dense deployment of LEO satellites, multi-satellite-enabled cooperative beamforming, which potentially enriches diversity and boosts received power, can be a promising solution.
@@ II-B1 Infrastructure Perspective
Owing to using higher frequency bands, LEO satellite communication systems can employ advanced infrastructures such as phased array antennas and hybrid beamforming structures to enhance communication performance, which also benefit positioning applications significantly. For instance, these infrastructures enable the acquisition of the angle-of-departure (AoD) and angle-of-arrival (AoA) observations, which are unavailable in GNSS that uses single-antenna transmitters and receivers. These angle observations are closely related to the UE position and can thus be explored to boost positioning performance. Moreover, deploying massive multiple-input multiple-output (MIMO) technology on LEO satellites enhances the accuracy of angle observations by providing significantly improved spatial direction resolution. Besides, in 5G/ sixth generation (6G) systems, the extensive terrestrial communication infrastructure can be leveraged and integrated with satellite-based positioning systems to enhance localization capabilities. Traditional GNSS applications already demonstrate the use of ground stations to assist high-precision positioning, as seen in real-time kinematic (RTK) positioning systems [12]. In emerging LEO satellite systems, theoretical studies have also shown that integrating ground-based technologies, such as reconfigurable intelligent surface (RIS), have the potential to significantly enhance positioning and tracking accuracy, particularly in complex suburban and urban environments [13].
@@ II-B2 Network Coordination Perspective
In addition to the advantages provided by new communication infrastructures, efficient inter-satellite and satellite-to-ground communication can also significantly enhance positioning performance by enabling cooperation across the network. For example, multiple LEO satellites equipped with MIMO antenna arrays can achieve superior performance through coordinated beamforming, as opposed to independent beamforming schemes. However, this coordination is only feasible with efficient inter-satellite communication links. Furthermore, it is preferable for the signals from each satellite to be orthogonal (in time or frequency), to mitigate mutual interference and ensure accurate positioning, which requires coordination between satellites. Therefore, information sharing between different positioning anchors (satellites or BSs), enabled by efficient communication links, is the key to enhancing the positioning performance of the entire system. However, this improvement comes at the cost of increased system complexity.
@@ II-B3 Other Benefits
LEO communication technologies enable satellites to synchronize their clocks with ground-based atomic clocks via signals, ensuring highly accurate time data for position calculations. Additionally, LEO communication technologies facilitate the implementation of fault-tolerant systems capable of automatically switching control from failed nodes to operational ones and rerouting data in the event of satellite or ground station failures. This redundancy is essential for maintaining uninterrupted positioning services. Furthermore, the two-way communication capability of LEO systems enhances dynamic and responsive positioning services, providing an improvement over systems lacking a return channel such as GNSS.
@@ III-A Simulation Setup
Considering a downlink scenario, we simulate LEO satellites at an altitude of 400 km, each with a transmit power of 60 dBm and a 240 MHz bandwidth, operating at 28 GHz. The ground UE is modeled as a single-antenna receiver with a noise power spectral density (PSD) of -174 dBm/Hz. The wireless channels are generated using the Rician fading model, with a Rician factor of 3 by assuming the user is in a rural environment [13, Table II]. By default, the LEO satellites are equipped with a $20\times 20$ antenna array with half-wavelength spacing, oriented towards the Earth. For simplicity, we employ an analog beamforming schemes.
@@ IV-A Doppler Shift
The strong Doppler effect on LEO satellite systems is a double-edged sword. On the one hand, the Doppler shift provides valuable information about the relative velocity between users and satellites, which is beneficial to positioning. On the other hand, this effect can be harmful to communication. For instance, a LEO satellite at an altitude of 800 km and operating at 30 GHz can experience a Doppler frequency shift variation rate of up to approximately 6 kHz/s [9]. This poses challenges for the phase-locked loop (PLL) used to mitigate Doppler shifts, as it must be rapidly adjusted to track the frequency shift beyond the initial offset caused by the rapid time variation. Thus, advanced algorithm and hardware are required to balance the utilization of the Doppler effect for accurate positioning and the mitigation of its negative effects in communication.
@@ IV-C Satellite Orbit Errors
The orbit information of LEO satellites is typically not accurately known due to several factors. Perturbing forces, such as atmospheric drag and solar radiation pressure, are primary contributors to these inaccuracies. Moreover, the process of parametrizing orbital data into two-line element (TLE) files introduces errors due to the constraints of the limited format. Combined with the use of simplified models for orbit propagation, these errors further degrade the accuracy of orbit predictions [5]. These factors introduce orbital errors, causing mismatches in the satellites’ position and orientation information, which ultimately degrade positioning and communication performance. Fig. 5 presents an evaluation of the positioning RMSE versus channel observation accuracy (using channel delay as a representative) under different levels of satellite position mismatch. The results indicate that such mismatches can lead to performance saturation, limiting the achievable accuracy of the system. A similar negative impact can be observed in communication performance, especially in an IPAC system.
@@ VI Acknowledgments
Fig. 1 was created by Ana Runte, a Scientific Illustrator at KAUST.